OpenStudy (anonymous):

if f(x)=2x^2+4 and g(x)=-x^3 then what is (f+g)(3) help

OpenStudy (anonymous):

\[f(x) = 2x^2 + 4 , g(x) = -x^3\]\[(f+g) = -x^3 + 2x^2 +4\]\[(f+g)(3) = -( 3)^3 +2(3)^2 + 4\]

OpenStudy (anonymous):

so it would be 24 ?

OpenStudy (anonymous):

how did you obtain 24 ?

OpenStudy (anonymous):

its : -27 + 18 + 4.. so -27 + 22 = -5 Ans: -5